write a quadratic equation with -3/4 and 4 as its roots, write the equation in form ax^2+bx+c=0, where a,b, and c are integers
Answer by Jk22(389)


We get : A(x-4)(x+3/4) as polynomial

= A (x^2 + 3x/4 -4x - 3)

= A (x^2 - 13x/4 - 3) = 0


hence A has to be a multiple of 4 (4m) in order the coefficient to be integers :

a = 4m, b = -13m, c = -12m, m integer (preferably non zero)